# Multilingual Brain Surgeon: Large Language Models Can be Compressed Leaving No Language Behind

### v1

@@ Abstract
[…]
@@ Appendix B MBS results tables
en 13.68 34.30 32.02 35.08 32.87
zh-Hans 23.70 72.92 101.79 62.39 59.28
fr 9.59 22.89 24.99 21.66 20.37
es 10.71 25.41 27.94 23.98 22.56
pt 10.97 27.47 34.99 25.82 24.27
ar 14.40 80.09 1.13E+14 41.36 47.57
vi 10.16 43.26 127.59 30.75 29.64
hi 10.96 44.10 1.20E+15 29.17 43.84
id 20.48 109.13 76038.83 64.40 79.85
bn 17.27 102.78 4.42E+23 56.65 121.11
ta 16.55 176.25 1.71E+07 64.40 209.90
te 18.10 355.98 6.96E+05 116.20 355.54
ur 13.26 97.65 3.62E+20 44.60 61.35
ne 27.22 555.21 1.27E+17 135.67 200.16
mr 23.07 503.51 4.25E+12 142.76 619.28
gu 21.52 327.11 2.17E+13 136.47 197.21
zh-Hant 21.84 137.00 266.49 61.30 58.56
sw 34.35 919.14 6011.36 357.63 336.80
yo 53.29 1.02E+03 3557.88 542.40 450.00
ig 39.16 728.01 1306.30 370.35 307.76
wikitext2 11.37 30.58 29.75 31.09 29.90
Average 20.08 257.99 2.11E+22 114.01 157.51
@@ Appendix C Monolingual pruning results
en 13.68 34.30 151% 49.17 259% 53.12 288% 38.50 181%
zh-Hans 23.70 72.92 208% 87.98 271% 126.56 434% 79.06 234%
fr 9.59 22.89 139% 32.46 238% 41.75 335% 24.91 160%
es 10.71 25.41 137% 38.29 258% 47.38 343% 28.18 163%
pt 10.97 27.47 150% 42.04 283% 57.68 426% 30.47 178%
ar 14.40 80.09 456% 79.26 451% 66.66 363% 55.03 282%
vi 10.16 43.26 326% 39.63 290% 150.26 1379% 38.05 275%
hi 10.96 44.10 303% 50.54 361% 32.37 195% 36.25 231%
id 20.48 109.13 433% 108.26 429% 952.66 4551% 60.77 197%
bn 17.27 102.78 495% 141.25 718% 68.01 294% 83.81 385%
ta 16.55 176.25 965% 147.99 794% 60.52 266% 163.29 887%
te 18.10 355.98 1866% 260.31 1338% 95.77 429% 445.67 2362%
ur 13.26 97.65 636% 143.11 979% 49.14 271% 63.39 378%
ne 27.22 555.21 1940% 320.61 1078% 143.19 426% 237.41 772%
mr 23.07 503.51 2083% 290.52 1159% 140.02 507% 220.62 856%
gu 21.52 327.11 1420% 215.13 900% 124.55 479% 286.66 1232%
zh-Hant 21.84 137.00 527% 156.15 615% 227.56 942% 167.85 668%
sw 34.35 919.14 2576% 419.52 1121% 897.07 2511% 465.66 1256%
yo 53.29 1024.94 1823% 529.35 893% 711.20 1235% 573.17 976%
ig 39.16 728.01 1759% 285.28 629% 773.60 1876% 409.11 945%
wikitext2 11.37 30.58 169% 40.70 258% 46.07 305% 33.40 194%
Average 20.08 257.99 883.9% 165.60 634.4% 231.67 850.2% 168.63 610.0%
@@
en 14 32 43 48 40
zh-Hans 24 102 128 108 110
zh-Hant 22 266 162 227 279
fr 10 25 28 39 28
es 11 28 34 48 32
pt 11 35 37 54 35
ar 14 1.13E+14 8.10E+07 247 3.52E+07
vi 10 128 70 948 78
hi 11 1.20E+15 3.93E+10 128 2.98E+09
id 20 76039 7049 11397 62
bn 17 4.42E+23 2.16E+10 1788 6.12E+15
ta 17 17114940 361068 85 1.80E+08
te 18 696319 180476 75 2.40E+07
ur 13 3.62E+20 55283 807 1.55E+07
ne 27 1.27E+17 1.70E+09 636 6.43E+10
mr 23 4.25E+12 2.13E+08 1250 3.76E+11
gu 22 2.17E+13 3051 252 4.96E+06
sw 34 6011 467 1363 1502
yo 53 3558 342 882 1463
ig 39 1306 205 990 753
wikitext2 11 30 36 41 34
Average 20 2.11E+22 2.99E+09 1020 2.92E+14

### v2

@@ Abstract
Hongchuan Zeng1, Hongshen Xu1, Lu Chen1,2†, Kai Yu1,2† 1X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China 2Suzhou Laboratory, Suzhou, China {charlie68, xuhongshen, chenlusz, kai.yu}@sjtu.edu.cn
@@ Appendix B MBS results tables
en 25.9744 34.30 32.02 35.08 32.87
zh-Hans 46.1386 72.92 101.79 62.39 59.28
zh-Hant 43.7887 137.00 266.49 61.30 58.56
fr 16.3204 22.89 24.99 21.66 20.37
es 18.1103 25.41 27.94 23.98 22.56
pt 19.1249 27.47 34.99 25.82 24.27
ar 28.1339 80.09 1.13E+14 41.36 47.57
vi 20.7135 43.26 127.59 30.75 29.64
hi 20.8597 44.10 1.20E+15 29.17 43.84
id 42.0467 109.13 76038.83 64.40 79.85
bn 37.0531 102.78 4.42E+23 56.65 121.11
ta 36.9749 176.25 1.71E+07 64.40 209.90
te 54.3996 355.98 6.96E+05 116.20 355.54
ur 29.419 97.65 3.62E+20 44.60 61.35
ne 75.7999 555.21 1.27E+17 135.67 200.16
mr 72.4524 503.51 4.25E+12 142.76 619.28
gu 70.6476 327.11 2.17E+13 136.47 197.21
sw 154.7551 919.14 6011.36 357.63 336.80
yo 267.0194 1.02E+03 3557.88 542.40 450.00
ig 185.7405 728.01 1306.30 370.35 307.76
wikitext2 22.4121 30.58 29.75 31.09 29.90
Average 61.33 257.99 2.11E+22 114.01 157.51
@@ Appendix C Monolingual pruning results
en 25.97 34.30 32% 49.17 89% 53.12 105% 38.50 48%
zh-Hans 46.14 72.92 58% 87.98 91% 126.56 174% 79.06 71%
zh-Hant 43.79 137.00 213% 156.15 257% 227.56 420% 167.85 283%
fr 16.32 22.89 40% 32.46 99% 41.75 156% 24.91 53%
es 18.11 25.41 40% 38.29 111% 47.38 162% 28.18 56%
pt 19.12 27.47 44% 42.04 120% 57.68 202% 30.47 59%
ar 28.13 80.09 185% 79.26 182% 66.66 137% 55.03 96%
vi 20.71 43.26 109% 39.63 91% 150.26 625% 38.05 84%
hi 20.86 44.10 111% 50.54 142% 32.37 55% 36.25 74%
id 42.05 109.13 160% 108.26 157% 952.66 2166% 60.77 45%
bn 37.05 102.78 177% 141.25 281% 68.01 84% 83.81 126%
ta 36.97 176.25 377% 147.99 300% 60.52 64% 163.29 342%
te 54.40 355.98 554% 260.31 379% 95.77 76% 445.67 719%
ur 29.42 97.65 232% 143.11 386% 49.14 67% 63.39 115%
ne 75.80 555.21 632% 320.61 323% 143.19 89% 237.41 213%
mr 72.45 503.51 595% 290.52 301% 140.02 93% 220.62 205%
gu 70.65 327.11 363% 215.13 205% 124.55 76% 286.66 306%
sw 154.76 919.14 494% 419.52 171% 897.07 480% 465.66 201%
yo 267.02 1024.94 284% 529.35 98% 711.20 166% 573.17 115%
ig 185.74 728.01 292% 285.28 54% 773.60 316% 409.11 120%
wikitext2 22.41 30.58 36% 40.70 82% 46.07 106% 33.40 49%
Average 61.33 257.99 239.5% 165.60 186.6% 231.67 277.0% 168.63 160.9%
@@
en 25.97 32.02 43.04 48.17 39.53
zh-Hans 46.14 101.79 127.73 107.62 109.50
zh-Hant 43.79 266.49 161.94 227.24 279.39
fr 16.32 24.99 28.45 38.85 27.65
es 18.11 27.94 34.01 47.80 32.46
pt 19.12 34.99 36.88 54.17 34.60
ar 28.13 1.13E+14 8.10E+07 246.62 3.52E+07
vi 20.71 127.59 69.64 947.65 77.75
hi 20.86 1.20E+15 3.93E+10 128.37 2.98E+09
id 42.05 76038.83 7049.31 1.14E+04 62.01
bn 37.05 4.42E+23 2.16E+10 1787.97 6.12E+15
ta 36.97 1.71E+07 3.61E+05 84.92 1.80E+08
te 54.40 6.96E+05 1.80E+05 74.68 2.40E+07
ur 29.42 3.62E+20 5.53E+04 807.08 1.55E+07
ne 75.80 1.27E+17 1.70E+09 635.82 6.43E+10
mr 72.45 4.25E+12 2.13E+08 1249.71 3.76E+11
gu 70.65 2.17E+13 3050.59 252.15 4.96E+06
sw 154.76 6011.36 466.51 1363.16 1502.12
yo 267.02 3557.88 342.02 881.53 1462.66
ig 185.74 1306.30 204.65 989.57 753.18
wikitext2 22.41 29.75 36.37 40.72 34.33
Average 61.33 2.11E+22 2.99E+09 1020 2.92E+14